OpenStudy (gokuporter):

Which answer is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 4.12(x 7.89) = 18.12 a)12.6 b)12.5 c) 12.4 d) 12.2

OpenStudy (gokuporter):

d?

pooja195 (pooja195):

I dont think your problem is typed correctly

OpenStudy (gokuporter):

It says it like that

pooja195 (pooja195):

Can you screenshot it

OpenStudy (gokuporter):

ok

OpenStudy (gokuporter):

http://prntscr.com/8wimpz

OpenStudy (gokuporter):

@pooja195

pooja195 (pooja195):

You forgot to add the minus sign in the equation u gave.

OpenStudy (gokuporter):

oh i see sorry

pooja195 (pooja195):

4.12(x-7.89) = 18.12 Round to the nearest number means this: 4(x-8)=19 do all the work and u get x=25/2 which also means 12.5

OpenStudy (gokuporter):

ok thank you for explaining it to me

pooja195 (pooja195):

No problem.